A family has two cars. The first car has a fuel efficiency of 25 miles per gallon of gas and the second has a fuel efficiency of 40 miles per gallon of gas. During one particular week, the two cars went a combined total of 2100 miles, for a total gas consumption of 60 gallons. How many gallons were consumed by each of the two cars that week?
Answer by oscargut(2103)

Let x = fuel consumed by the first car
Let y = fuel consumed by the second car
Then:
x+y = 60
25x+40y = 2100
Then:
25x+40(60-x)=2100
-15x+2400 = 2100
15x = 300
Answer:
x = 20
y = 40

You can ask me more at: mthman@gmail.com
Thanks